OpenStudy (anonymous):

Integral of x/(x+2)^(1/2) from -1 to 2

OpenStudy (anonymous):

\[\int\limits_{2}^{-1}(x/\sqrt(x+2)\]

OpenStudy (anonymous):

take anti-derivative and solve

OpenStudy (anonymous):

a cud u help me

OpenStudy (anonymous):

for F(b)-F(a)

OpenStudy (anonymous):

integration by parts: u=x, du=dx, dv=(x+2)^(-1/2), v=2*(x+2)^(1/2) so, 2x*(x+2)^1/2 -2int((x+2)^(1/2)dx=....2/3